OpenStudy (anonymous):

what are the factors of the expression x^2+x-20

OpenStudy (wearehiding):

factor | x^2+x-20 (x-4) (x+5) and btw this is the History Section.

OpenStudy (anonymous):

how would you solve this problem step by step

OpenStudy (wearehiding):

x^2 - x - 20 ----(1) =x^2 - 5x+4x - 20 (splitting the middle term (-x) into two terms so that their sum is (-x) and their product is the product of the square term and the constant term.) =(x^2 - 5x)+ (4x - 20) =x(x-5)+4(x-5) =xp+4p where p= (x-5) = p(x+4) =(x-5)(x+4) This means (x-5) and (x+4) are factors

OpenStudy (anonymous):

i think you made a mistake

OpenStudy (anonymous):

@WeAreHiding

OpenStudy (wearehiding):

@e.mccormick

OpenStudy (anonymous):

x^2 - x - 20 (Note that x^2 means x squared). The coefficient of x (the second term in the equasion) is negative one. (Remember that anything times 1 equals itself, so -1x is the same thing as -x). We'll re-write the expression to make that clear: x^2 - 1x - 20 The general format of a factored equasion is (x + ?)(x + ?). We need to fill in the question marks. If you have to numbers, you want them to equal -20 (the third term) when multiplied and -1 (the coefficient of the second term) when added. So, what numbers, when multiplied, equal negative 20? Let's try a few combinations: 5 and -4 -5 and 4 10 and -2 2 and -10 Which one of these combinations, when added together, equals negative 1? Negative 5 and 4! Fill those into the question marks: The answer is: (x - 5 )(x + 4 ) You can check it by FOIL-ing it out (FOIL stands for First, Outer, Inner, Last. It refers to the order that you should multiply the terms in). First terms: x * x = x^2 Outer terms: x * 4 = 4x Inner: -5 * x = -5x Last: -5 * 4 = -20 So, we have x^2 + 4x -5x -20 Add like terms, and you get x^2 - 1x - 20 -- the original equasion, so you know that (x - 5)(x + 4) is the right answer!

OpenStudy (anonymous):

@WeAreHiding is correct

OpenStudy (anonymous):

So how would i write it

OpenStudy (anonymous):

Hun, we can't just give you the answer to copy. That goes against Open Study's policies. Open study tries to have student help each other learn. Review what I've sent you and try to understand it. If you need help I'll try to help, but you really need to learn this stuff. If you still do not know how to answer or show work, I suggest that you go and ask your teacher for help

OpenStudy (anonymous):

No, i understand but the thing is that i dont know how to write it to begin with

OpenStudy (anonymous):

mainly what I said but with the words turned into the equations Im describing